A 46.5 g ice cube, initially at 0°C, is dropped into a Styrofoam cup containing 384 g of water, initially at 21.6°C. What is the final temperature of the water, if no heat is transferred to the Styrofoam or the surroundings?
Answer #1

ice haat Iast by wate Heat gaine by At, m, L m, bet final fomperature be T At, = (T-0) c hen At= (21-6-7) mass of ice melteou6.Sg X334J9 +4g x 4.186 Jgx 384g X 4186丁g-2x(216-T)と 15531 94.649T 34720.3584 - 1607.424T 1802.073T= 19189-3584 lo 6485